OpenStudy (anonymous):

prove: \[\cos ^{2}x-\frac{ 1 }{ 2 }tanxsin(2x)=\cos(2x)\]

OpenStudy (solomonzelman):

If \(\color{#000000 }{ \displaystyle \cos^2x-\frac{1}{2}\tan x \sin(2x)=\cos(2x) }\) then, \(\color{#0000f0 }{ \displaystyle \frac{1}{2}\tan x\sin(2x) =\sin^2x}\) since \(\color{#000000 }{ \displaystyle \cos(2x)=\cos^2x-\sin^2x }\)

OpenStudy (solomonzelman):

So, if you prove \(\color{#0000f0 }{ \displaystyle \frac{1}{2}\tan x\sin(2x) =\sin^2x}\) then your hypothesis is true

OpenStudy (solomonzelman):

Note that \(\color{#000000 }{ \displaystyle \sin(2x) =2\sin x \cos x}\)

OpenStudy (solomonzelman):

From there should be really easy...

OpenStudy (anonymous):

yes i got it now and i have 1 more proof its : \[\frac{ cosxtanx+sinx }{ cosx }=2tanx\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{ax+b }{a}=\frac{ax }{a} +\frac{b}{a}=x+\frac{b}{a} }\)

OpenStudy (solomonzelman):

you would agree with this kind of statement, right?

OpenStudy (solomonzelman):

So, just separate the fraction (as I did with a,b,x), and simplify.

OpenStudy (solomonzelman):

Note that you know sin(x)/cos(x)=tan(x) And regardless of what Z is, Z+Z=2Z.

OpenStudy (anonymous):

yes i aree with that

OpenStudy (solomonzelman):

ok, then you should be able to simplify that, apply basic trig and you are good to go:)

OpenStudy (anonymous):

but when you said separate do you mean like this \[\frac{ cosx }{ cosx}+\frac{ tanx }{ cosx }+\frac{ sinx }{ cosx }\]

OpenStudy (solomonzelman):

absolutely not

OpenStudy (anonymous):

well with the cos and tan together sorry

OpenStudy (solomonzelman):

yes, they are together, so you can just replace the first + you put, with ×

OpenStudy (anonymous):

yea and tangent is just sinx/cosx

OpenStudy (solomonzelman):

yes, but I would ay it the other way aroundfor the purpose of your identity

OpenStudy (anonymous):

but on one side its \[\frac{ cosxtanx }{ cosx}+\frac{ sinx }{ cosx }\]

OpenStudy (solomonzelman):

ok, so the first time you cancel the cos(x) on top and bottom, and you get? sin(x)/cos(x) is same thing as? then add tangents. look at the right side

OpenStudy (anonymous):

ohhhhhhhh!! once it cancels you just get 2 ttangents basicaly

OpenStudy (solomonzelman):

yes, that is exactly right!

OpenStudy (anonymous):

thanks so much

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

hey the first proof we did i dont really understand

OpenStudy (anonymous):

\[\cos ^{2}x-\frac{ 1 }{ 2 }tanxsin(2x)=\cos(2x)\]

OpenStudy (solomonzelman):

Ok, there is a rule: \(\color{#000000 }{ \displaystyle\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b) }\)

OpenStudy (solomonzelman):

We can prove it geometrically if you want, or if you want to take that as given, then that is fine too

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

And also, \(\color{#000000 }{ \displaystyle\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b) }\)

OpenStudy (anonymous):

ill take the second option lol

OpenStudy (solomonzelman):

no no, the first one is the one we apply, I am just letting you know of it

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b) }\) and what happens when a and b are the same number, let say that both are x.

OpenStudy (anonymous):

2x

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle\cos(2x)=\cos(x+x)=\cos(x)\cos(x)-\sin(x)\sin(x) }\)

OpenStudy (solomonzelman):

and that is same as \(\color{#000000 }{ \displaystyle\cos(2x)=\cos^2(x)-\sin^2(x) }\)

OpenStudy (anonymous):

yes so can you say sin(2x)=2sinxcosx

OpenStudy (solomonzelman):

So, \(\color{#000000 }{ \displaystyle \cos^2(x)-\frac{1}{2}\sin(2x)\tan(x)=\cos(2x)}\) \(\color{#000000 }{ \displaystyle \cos^2(x)-\frac{1}{2}\sin(2x)\tan(x)=\cos^2(x)-\sin^2(x)}\) \(\color{#000000 }{ \displaystyle \frac{1}{2}\sin(2x)\tan(x)=\sin^2(x)}\)

OpenStudy (solomonzelman):

yes, and then you apply \(\sin(2x)=2\sin(x)\cos(x)\)

OpenStudy (solomonzelman):

And that rule comes from \(\color{#000000 }{ \displaystyle \sin(a+b)=\sin(a)\cos(b)+\sin(b)\cos(a)}\) (after setting a and b =x)

OpenStudy (anonymous):

yes but how to conclude that that equals cos(2x)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \cos^2(x)-\frac{1}{2}\sin(2x)\tan(x)=\cos^2(x)-\sin^2(x)}\) \(\color{#000000 }{ \displaystyle \cos^2(x)-\frac{1}{2}\cdot (2\sin(x)\cos(x))\cdot \tan(x)=\cos^2(x)-\sin^2(x)}\) \(\color{#000000 }{ \displaystyle \cos^2(x)-\frac{1}{2}\cdot (2\sin(x)\cos(x))\cdot \frac{\sin(x)}{\cos(x)}=\cos^2(x)-\sin^2(x)}\)

OpenStudy (solomonzelman):

simplify that and boom boom :)

OpenStudy (anonymous):

i got it now

OpenStudy (solomonzelman):

cool